o.k., one more for today, thank you everyone who helped me!
A temperature of 41 degrees F is equivalent to 278.2 degrees K, while a temp. of 86 degrees F is equivalent to 303.2 degrees K. Find the converting Kelvin temp. to Fahrenheit. What is the Fahrenheit temp. when it is 180 degrees K?
Word problems stump me....

Answer by Edwin McCravy(20056)

o.k., one more for today, thank you everyone who helped me! 
A temperature of 41 degrees F is equivalent to 278.2 degrees K,
while a temp. of 86 degrees F is equivalent to 303.2 degrees K. 
Find the converting Kelvin temp. to Fahrenheit. What is the 
Fahrenheit temp. when it is 180 degrees K? 
Word problems stump me....

===========================================================

This is the problem of finding the equation of the line that
passes through two points.

Let x = Kelvin temperature
Let y = Fahrenheit temperature

>>...41 degrees F is equivalent to 278.2 degrees K...<<

Translation: when x = 278.2, y = 41 OR

The line passes through the point (278.2, 41)

>>...56 degrees F is equivalent to 303.2 degrees K...<<

Translation: when x = 303.2, y = 86 OR

The line passes through the point (303.2, 86)

Use the slope formula

     y2 - y1        86 - 41       9   
m = --------- = --------------- = -- = 1.8
     x2 - x1     303.2 - 278.2    5     

Use the point-slope formula

y - y1 = m(x - x1)

y - 41 = 1.8(x - 278.2)

y - 41 = 1.8x - 500.76

     y = 1.8x - 459.76

We can change the x to K and the y to F:

     F = 1.5K - 459.76

>>...What is the Fahrenheit temp. when it is 180 degrees K?...<<

     F = 1.5(180) - 459.76

     F = -135.76°

Edwin